OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch radius A = sq. in.

OpenStudy (anonymous):



OpenStudy (anonymous):

can you do it from here? that small triangle is a 30-60-90 triangle...

OpenStudy (anonymous):

A = 1/2aP A = 1/2(3)(P) A = 3/2(P) To find the perimeter, use the 30-60-90 ratio, find the side length of the side opposite the 60º angle and multiply it by 6.

OpenStudy (anonymous):

A = 3/2(18√3) A = 27√3

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):